# AttackSeqBench: Benchmarking Large Language Models in Analyzing Attack Sequences within Cyber Threat Intelligence

### v2

@@ B.1 TTP Temporal Position Analysis
We summarize our findings into three key points: (1) Comparing the performance of three benchmark tasks, the overall highest standard deviation occurs within the AttackSeq-Tactic task for all LLMs and LRMs, suggesting that the LLMs’ and LRMs’ performance has the largest variance in this taskand further highlighting the challenge of inferring tactics within attack sequences; (2) In contrast, the standard deviation for AttackSeq-Procedure is the lowest for all LLMs, suggesting more consistency in the LLMs’ and LRMs’ performance among the different attack sequence stages; (3) GPT-4o consistently outperforms other LLMs and LRMs across most benchmark tasks, particularly in the Exploitation Phase and the Stealth Expansion Phase. However, the remaining LLMs achieve comparable performance in the remaining two phases, especially in tactics at the start and end of the attack sequences (e.g., Reconnaissance and Impact). This suggests that TTPs within rarely observed tactics are more challenging for the LLM and LRMs to interpret.
B.2 Case Study
Table 7: Error analysis of GPT-4o in AttackSeq-Technique, where GPT-4o incorrectly associates technique T1574-Hijack Execution Flow with the Execution tactic.
@@ B.2 Case Study
Table 8: Performance comparison between three embedding models (abbreviate as Emb. M.).
#Emb. M. BGE OPENAI ATT&CK-BERT
@@
R1 (Llama-8B) 0.4905 0.5740 0.5226 0.4932 0.5696 0.5150 0.4651 0.5804 0.5104
GPT o3-mini 0.5115 0.5853 0.6474 0.5192 0.5827 0.6474 0.5245 0.5874 0.6414
B.3 Impact of Embedding Models within RAG-empowered Setting
The semantics within CTI reports contain a large volume of domain-specific technical terminologies, where the accuracy of retrievers in identifying the most relevant tactics, techniques and procedures critically influences LLMs’ performance in RAG-empowered settings. To further examine this issue and mitigate the potential knowledge bias introduced by BGE-EN-ICL, we incorporate two additional embedding models into our RAG-empowered setting, namely OpenAI’s text-embedding-3-large (OpenAI, 2024b) and a domain-adapted ATT&CK BERT  (Abdeen et al., 2023b) fine-tuned on the specific cybersecurity data. We conduct performance comparisons between two representative LLMs and two LRMs based on these above embedding models in Table 8. We observe that the three embedding models exhibit comparable performance across different benchmark tasks and settings, with only marginal differences. Among them, BGE-EN-ICL proves to be the most cost-efficient, generalizable, and effective choice, and thus we primarily report model performance based on this embedding throughout the paper. Notably, although ATT&CK-BERT contains far fewer parameters than BGE-EN-ICL (110M vs. 7B), it achieves comparable performance, underscoring the importance of injecting domain-specific security knowledge into LLMs and pointing to a promising direction for future work.
@@ Appendix C Prompt Templates
C.1 Question Generation Prompt Templates
We utilize few-shot prompt templates for question generation of each of our benchmark tasks (i.e., AttackSeq-Tactic, AttackSeq-Technique, AttackSeq-Procedure-Yes, AttackSeq-Procedure-No), the corresponding prompts are shown in Box C.1, Box C.1, Box C.1, Box C.1 respectively.
Box 1: Few-shot prompt for question generation for AttackSeq-Tactic. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]:
@@ 3. Provide the Question-Answer Pair:
Box 2: Few-shot prompt for question generation for AttackSeq-Technique. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A MITRE ATT&CK Tactic present in the CTI report. 3. A MITRE ATT&CK Technique present in the CTI report. [End of Inputs]
@@
Box 3: Few-shot prompt for question generation for AttackSeq-Procedure-Yes. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive four parts: 1. A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A MITRE ATT&CK Tactic present in the CTI report. 3. A MITRE ATT&CK Technique present in the CTI report. 4. 4. A list of Procedures present in the CTI report, where each procedure is represented as a (Subject, Relation, Object) triplet. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, the question should focus on inferring the given list of procedures based on the given MITRE ATT&CK tactic and technique. The answer to the question is ”Yes”, indicating that the given list of procedures is likely to occur in the attack sequence. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is ”Yes”. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, is it likely that the Ke3chang malware will run commands on the command-line interface before achieving persistence by adding a Run key? Answer: Yes - Question: Is it likely that Axiom will acquire dynamic DNS services for use in the targeting of intended victims before gaining initial access to the victim’s network using SQL injection? Answer: Yes - Question: Is Ke3chang likely to perform frequent and scheduled data exfiltration from compromised networks after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: Yes - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, is APT41 likely to use Windows services to execute DUSTPAN before using Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: Yes [End of Examples]
@@
Box 4: Few-shot prompt for question generation for AttackSeq-Procedure-No. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]:
@@
1. A Reference Question-Answer Pair that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always ”Yes”. 2. A Reference MITRE TTP that is NOT supported in the CTI report. [End of Inputs]
Your task is to generate two questions based on the given reference question, such that attack sequence described in the question is modified and the correct answer to the two questions is ”No”. The definitions of the two questions are as follows: 1. Question 1 should negate the ”before” and/or ”after” clauses of the reference question, such that the attack sequence contradicts the original sequence of TTPs in the reference question. 2. Question 2 should replace the main procedures in the reference question with the provided reference MITRE TTP, such that the replaced main procedures are not found in the CTI report. [Steps]:
@@
1. Analyze the Reference Question-Answer Pair: - Identify and outline the attack sequence in the order presented in the reference Question-Answer Pair. 2. Construct the Questions: - Design two questions that modify the attack sequence described in the reference question. - Ensure that the answer to the questions is ”No”. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the examples below for guidance. [Examples]:
@@ 3. Provide the Question-Answer Pairs:
C.2 Dataset Refinement Prompt Templates
The prompt templates to filter based on the Answerability criteria is in Box C.2, while the feedback and refinement prompts are in Box C.2 and Box C.2 respectively.
Box 5: Prompt template for verifying Answerability during Self-Refinement. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. TTP Description: A reference description of the correct answer corresponding to the question. 3. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. [End of Inputs] Your task is to evaluate the answerability of the given question using the provided information in the CTI Outline and TTP description. We define answerability based on three factors below: 1. The correct answer must be supported by the CTI outline. 2. The correct answer must clearly stand out as the best answer choice to the question based on the CTI outline. 3. Suppose the masked_tactic paragraph is removed from the CTI outline, the correct answer must be deducible from the answer choices by using the information provided in remaining tactics of the CTI outline and TTP description. You may also refer to your external cybersecurity knowledge to determine if the correct answer is deducible. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Match the correct answer with the provided TTP description. - Determine step-by-step if the question is answerable based on the definition above. 3. Output evaluation result: - Output one of the following: - “A”: Indicates that the question is answerable. - “B”: Indicates that the question is not answerable. - “C”: Indicates that you do not know/cannot determine if the question is answerable. - Please also include a short and concise explanation of your evaluation result. - Please follow the output format: “Explanation: <insert explanation here> Evaluation Result: <insert letter here>.” [End of Steps]
@@ C.2 Dataset Refinement Prompt Templates
Box 6: Prompt template for assessing question quality based on the evaluation criteria. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the Tactics, Techniques and Procedures (TTPs) in MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. TTP Description: A reference description of the correct answer to the question. 3. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI outline, along with one correct answer and distractors among the answer choices. [End of Inputs]
Your task is to evaluate the QA pair and provide your feedback for each of the criteria defined below: [Evaluation Criteria]:
@@
Box 7: Prompt template for question refinement. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. Feedback Results: A list of feedback scores and explanations for each desired criterion of the question defined below. [End of Inputs] Your task is to iteratively refine the quality of the given question based on the feedback provided in the Feedback Results. [Evaluation Criteria]: Please refer to the definition of each feedback criterion: 1. Clarity: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. Logical: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. Relevance: Does the question directly relate to the CTI outline? 4. Consistency: Does the question align with the provided TTP Description? 5. Answer Consistency: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. 2. Analyze the Question with Answer Choices: - Read the question and the provided answer choices carefully. 3. Analyze the Feedback Results: - Based on the feedback given in each criterion, refine the question to improve the each aspect. - Please ensure that the correct answer to the refined question is the same as the original question. - Please also ensure that the question avoids hinting at the correct answer. 4. Output the Refined Question: - Please follow the output format: “Refined Question: <Your refined question here>.” [End of Steps] Following the steps above, please refine the question based on the Feedback Results and CTI Outline provided below. Please only output the refined question.
C.3 Automatic Evaluation Prompt Templates
We utilize the definitions in the evaluation criteria in Table LABEL:tab.annotation_instructions to create prompts. We show an example prompt template for evaluating the Logical aspect of the question shown in Box C.3.
@@ C.3 Automatic Evaluation Prompt Templates
Box 8: Prompt template for evaluating the Logical aspect. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. Description to Correct Answer: The description of the correct answer from the MITRE ATT&CK framework. [End of Inputs] Your task is to rate the question on one metric below. [Definition]

### v3

@@ B.1. TTP Temporal Position Analysis
To better understand the LLMs’ capabilities in attack sequence understanding, we conduct fine-grained analysis based on each stage within the attack sequences of MITRE ATT&CK®. We illustrate the performance of two LLMs and two LRMs across all benchmark tasks in the Regular setting in Figure 9 and show the corresponding values, the mean and standard deviation (SD) of these LLMs and LRMs on each tactic and benchmark task in Table 6. We identify four overachieving attack phases to categorize the ATT&CK tactics in attack sequences: (1) Initial Intrusion Phase; (2) Exploitation Phase; (3) Stealth Expansion Phase; (4) Objective Orchestration Phase. It is worth noting that our categorization follows Tactics, as each Technique and Procedure in MITRE ATT&CK® is uniquely mapped to a specific Tactic within a given attack sequence.
We summarize our findings into three key points: (1) Comparing the performance of three benchmark tasks, the overall highest standard deviation occurs within the AttackSeqBench-Tactic task for all LLMs and LRMs, suggesting that the LLMs’ and LRMs’ performance has the largest variance in this task and further highlighting the challenge of inferring tactics within attack sequences; (2) In contrast, the standard deviation for AttackSeqBench-Procedure is the lowest for all LLMs, suggesting more consistency in the LLMs’ and LRMs’ performance among the different attack sequence stages; (3) GPT-4o consistently outperforms other LLMs and LRMs across most benchmark tasks, particularly in the Exploitation Phase and the Stealth Expansion Phase. However, the remaining LLMs achieve comparable performance in the remaining two phases, especially in tactics at the start and end of the attack sequences (e.g., Reconnaissance and Impact). This suggests that TTPs within rarely observed tactics are more challenging for the LLM and LRMs to interpret.
Table 7. Error analysis of GPT-4o in AttackSeqBench-Technique, where GPT-4o incorrectly associates technique T1574-Hijack Execution Flow with the Execution tactic.
@@
B.2. Case Study
Table 8. Performance comparison between three embedding models (abbreviate as Emb. M.).
#Emb. M. BGE-EN-ICL OPENAI ATT&CK-BERT
@@ B.2. Case Study
R1 (LLaMA3.1-8B) 0.4905 0.5740 0.5226 0.4932 0.5696 0.5150 0.4651 0.5804 0.5104
GPT-o3-mini 0.5115 0.5853 0.6474 0.5192 0.5827 0.6474 0.5245 0.5874 0.6414
B.3. Impact of Embedding Models within RAG-empowered Setting
The semantics within CTI reports contain a large volume of domain-specific technical terminologies, where the accuracy of retrievers in identifying the most relevant tactics, techniques and procedures critically influences LLMs’ performance in RAG-empowered settings. To further examine this issue and mitigate the potential knowledge bias introduced by BGE-EN-ICL, we incorporate two additional embedding models into our RAG-empowered setting, namely OpenAI’s text-embedding-3-large (OpenAI, 2024b) and a domain-adapted ATT&CK BERT  (Abdeen et al., 2023a) fine-tuned on the specific cybersecurity data. We conduct performance comparisons between two representative LLMs and two LRMs based on these above embedding models in Table 8. We observe that the three embedding models exhibit comparable performance across different benchmark tasks and settings, with only marginal differences. Among them, BGE-EN-ICL proves to be the most cost-efficient, generalizable, and effective choice, and thus we primarily report model performance based on this embedding throughout the paper. Notably, although ATT&CK-BERT contains far fewer parameters than BGE-EN-ICL (110M vs. 7B), it achieves comparable performance, underscoring the importance of injecting domain-specific security knowledge into LLMs and pointing to a promising direction for future work.
@@ Appendix C Prompt Templates
C.1. Question Generation Prompt Templates
We utilize few-shot prompt templates for question generation of each of our benchmark tasks (i.e., AttackSeqBench-Tactic, AttackSeqBench-Technique, AttackSeqBench-Procedure-Yes, AttackSeqBench- Procedure-No), the corresponding prompts are shown in Box C.4, Box C.4, Box C.4, Box C.4 respectively.
C.2. Dataset Refinement Prompt Templates
The prompt templates to filter based on the Answerability criteria is in Box C.4, while the feedback and refinement prompts are in Box C.4 and Box C.4 respectively.
C.3. Automatic Evaluation Prompt Templates
We utilize the definitions in the evaluation criteria in Table 5 to create prompts. We show an example prompt template for evaluating the Logical aspect of the question shown in Box C.4.
C.4. Answering Prompt Templates
Box 1: Few-shot prompt for question generation for AttackSeqBench-Tactic. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]:
@@ 3. Provide the Question-Answer Pair:
Box 2: Few-shot prompt for question generation for AttackSeqBench-Technique. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A MITRE ATT&CK Tactic present in the CTI report. 3. A MITRE ATT&CK Technique present in the CTI report. [End of Inputs]
@@
Box 3: Few-shot prompt for question generation for AttackSeqBench-Procedure-Yes. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive four parts: 1. A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A MITRE ATT&CK Tactic present in the CTI report. 3. A MITRE ATT&CK Technique present in the CTI report. 4. 4. A list of Procedures present in the CTI report, where each procedure is represented as a (Subject, Relation, Object) triplet. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, the question should focus on inferring the given list of procedures based on the given MITRE ATT&CK tactic and technique. The answer to the question is "Yes", indicating that the given list of procedures is likely to occur in the attack sequence. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is "Yes". - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, is it likely that the Ke3chang malware will run commands on the command-line interface before achieving persistence by adding a Run key? Answer: Yes - Question: Is it likely that Axiom will acquire dynamic DNS services for use in the targeting of intended victims before gaining initial access to the victim’s network using SQL injection? Answer: Yes - Question: Is Ke3chang likely to perform frequent and scheduled data exfiltration from compromised networks after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: Yes - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, is APT41 likely to use Windows services to execute DUSTPAN before using Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: Yes [End of Examples]
@@
Box 4: Few-shot prompt for question generation for AttackSeqBench-Procedure-No. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]:
@@
1. A Reference Question-Answer Pair that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always "Yes". 2. A Reference MITRE TTP that is NOT supported in the CTI report. [End of Inputs]
Your task is to generate two questions based on the given reference question, such that attack sequence described in the question is modified and the correct answer to the two questions is "No". The definitions of the two questions are as follows: 1. Question 1 should negate the "before" and/or "after" clauses of the reference question, such that the attack sequence contradicts the original sequence of TTPs in the reference question. 2. Question 2 should replace the main procedures in the reference question with the provided reference MITRE TTP, such that the replaced main procedures are not found in the CTI report. [Steps]:
@@
1. Analyze the Reference Question-Answer Pair: - Identify and outline the attack sequence in the order presented in the reference Question-Answer Pair. 2. Construct the Questions: - Design two questions that modify the attack sequence described in the reference question. - Ensure that the answer to the questions is "No". - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the examples below for guidance. [Examples]:
@@ 3. Provide the Question-Answer Pairs:
Box 5: Prompt template for verifying Answerability via Self-Refinement. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under "Others" provides background on the threat actor but is secondary. 2. TTP Description: A reference description of the correct answer corresponding to the question. 3. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. [End of Inputs] Your task is to evaluate the answerability of the given question using the provided information in the CTI Outline and TTP description. We define answerability based on three factors below: 1. The correct answer must be supported by the CTI outline. 2. The correct answer must clearly stand out as the best answer choice to the question based on the CTI outline. 3. Suppose the masked_tactic paragraph is removed from the CTI outline, the correct answer must be deducible from the answer choices by using the information provided in remaining tactics of the CTI outline and TTP description. You may also refer to your external cybersecurity knowledge to determine if the correct answer is deducible. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Match the correct answer with the provided TTP description. - Determine step-by-step if the question is answerable based on the definition above. 3. Output evaluation result: - Output one of the following: - “A”: Indicates that the question is answerable. - “B”: Indicates that the question is not answerable. - “C”: Indicates that you do not know/cannot determine if the question is answerable. - Please also include a short and concise explanation of your evaluation result. - Please follow the output format: “Explanation: <insert explanation here> Evaluation Result: <insert letter here>.” [End of Steps]
@@
Box 6: Prompt template for assessing question quality based on the evaluation criteria. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the Tactics, Techniques and Procedures (TTPs) in MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under "Others" provides background on the threat actor but is secondary. 2. TTP Description: A reference description of the correct answer to the question. 3. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI outline, along with one correct answer and distractors among the answer choices. [End of Inputs]
Your task is to evaluate the Q&A pair and provide your feedback for each of the criteria defined below: [Evaluation Criteria]:
@@
Box 7: Prompt template for question refinement. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under "Others" provides background on the threat actor but is secondary. 2. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. Feedback Results: A list of feedback scores and explanations for each desired criterion of the question defined below. [End of Inputs] Your task is to iteratively refine the quality of the given question based on the feedback provided in the Feedback Results. [Evaluation Criteria]: Please refer to the definition of each feedback criterion: 1. Clarity: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. Logical: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. Relevance: Does the question directly relate to the CTI outline? 4. Consistency: Does the question align with the provided TTP Description? 5. Answer Consistency: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. 2. Analyze the Question with Answer Choices: - Read the question and the provided answer choices carefully. 3. Analyze the Feedback Results: - Based on the feedback given in each criterion, refine the question to improve the each aspect. - Please ensure that the correct answer to the refined question is the same as the original question. - Please also ensure that the question avoids hinting at the correct answer. 4. Output the Refined Question: - Please follow the output format: “Refined Question: <Your refined question here>.” [End of Steps] Following the steps above, please refine the question based on the Feedback Results and CTI Outline provided below. Please only output the refined question.
@@
Box 8: Prompt template for evaluating the Logical aspect. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under "Others" provides background on the threat actor but is secondary. 2. Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. Description to Correct Answer: The description of the correct answer from the MITRE ATT&CK framework. [End of Inputs] Your task is to rate the question on one metric below. [Definition]